Medicine is packed in boxes, each weighing 4 kg 500g. How many such boxes can be loaded in a van which cannot carry beyond 800 kg?

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Weight of one box = 4 kg 500 g = 4×1000 + 500
= 4500 g
Maximum weight carried by the van = 800 kg = 800×1000
= 800000 g
Hence, number of boxes that can be loaded in the van = 800000 / 4500 = 177 boxes

Thanks